OpenStudy (angeleyes108):

Please Help!! Fan & Medal: 50/? = 230/1380 (Equation written better below)

OpenStudy (angeleyes108):

\[\frac{ 50 }{ ? } = \frac{ 230 }{ 1380 }\]

OpenStudy (campbell_st):

start by cross multiplying \[50 \times 1380 = ? \times 230\] now divide both sides by 230 \[\frac{50 \times 1380}{230} = ?\] hope it helps

OpenStudy (angeleyes108):

Oh that helps a lot! Thank you!